OpenStudy (anonymous):

4 C3H5N3O9---> 6 N2+ 12 CO2 + 10 H2O + O2 This reaction generates a large amount of heat and many gaseous products. It is the sudden formation of these gases, together with a rapid expansion, that produces the explosion. a) what is the maximum amount of O2 in grams that can be obtained from 2.00x10^2 of nitroglycerin? b) calculate the percent yield in this reaction if the same amount of O2 generated is found to be 6.55g.

OpenStudy (aaronq):

convert the mass of nitroglygerin you have to moles. then build a ratio Using the stoichiometric coefficients to find moles produced. -------------------------------------------------------------- e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients, \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) so if you have moles of B= 2, how many moles of C did you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_c=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. b) What you found in part a) was the "theoretical yield". They're telling you th actual yield is 6.55g of \(O_2\). So use this: Percent yield = \(\dfrac{actual\;yield}{theoretical \;yield}*100\%\)